OpenStudy (lgbasallote):

\(\LARGE \int_{0}^{\frac{\pi}{4}} \sin^2 (4x) \cos^2 (2x) dx\) Solve by Wallis Formula anyone wanna show me a demo? :DDD like..say..uhmm... @Mimi_x3 ? hehe

OpenStudy (lgbasallote):

i just need to manipulate it to the form \(\LARGE \int_{0}^{\frac{\pi}{2}} \sin^m u \cos^n u du\) when m,n \(\ge\) 0 u = 0 \(\large u_{2} = \frac{\pi}{2}\) i knoow the rest :D

OpenStudy (aravindg):

:)

OpenStudy (lgbasallote):

spill it if you know :p

OpenStudy (aravindg):

here we go

OpenStudy (lgbasallote):

where'd we go?

OpenStudy (aravindg):

well ...cant we do this without wallis formula?

OpenStudy (lgbasallote):

hmm i think i can :P but i need to learn the wallis formula...

OpenStudy (aravindg):

if its without wallis formula i can try ..nt learned that yet :P

OpenStudy (lgbasallote):

well i just need help manipulating it to that form...just need to see how

OpenStudy (aravindg):

i gt it!!

OpenStudy (aravindg):

use sin 2x formula

OpenStudy (aravindg):